OpenStudy (kkbrookly):

Can someone please check my answers??? Given tan theta=-4/5 and cos theta > 0, find the exact values of sin theta and sec theta. Sketch the problem situation. Label the sketch with the given information. Determine the needed information.

OpenStudy (anonymous):

make a triangle

OpenStudy (kkbrookly):

I have -0.63 radians for sin theta and 1.28 for sec theta

OpenStudy (anonymous):



OpenStudy (anonymous):

there is a picture of an angle whose tangent is \(\frac{4}{5}\) find the hypotenuse via pythagoras then you can find any trig ratio you like just be careful as to what quadrant you are in

OpenStudy (kkbrookly):

Alright, thank you! Are my answers correct though?

OpenStudy (anonymous):

you are not being asked for the measure of the angle, just for the other trig ratios they probably want them written in radical form, not as decimals

OpenStudy (kkbrookly):

Oh alright. That makes sense.

OpenStudy (anonymous):

also \(\sin(\theta)\) is a number, not an angle

OpenStudy (kkbrookly):

How do I convert my decimals into radical form?

OpenStudy (anonymous):

you don't find the hypotenuse of the triangle above, then take the trig ratios you want

OpenStudy (kkbrookly):

Thank you!